OpenStudy (anonymous):

The cost of a taxi ride is given byC = rd + a, where r is the rate per one-fifth mile, d is the trip distance in terms of the number of one-fifth-mile units in the trip, and a is the flag-drop fee (an automatic charge created when the meter is started). Solve the equation for the mileage rate r. Pleaseand be detailed.

OpenStudy (anonymous):

@Rachel98 This is obviously a word problem, but don't let the words confuse you... they actually don't matter much at all here. The question gives you one expression C = rd + a which is an equation "solved for" C, because C is "all alone" on the left side of the equals sign. It asks you to solve for "r" which means you need to get r all alone on one side of the equals sign. You can change the equation however you need, but the BIG rule is that anything you do on one side of the equation, you MUST do on the other. C = rd + a C -a = rd + a - a <<--- subtract "a" from both sides C-a = rd <<---- simplify, since a-a = 0 (C-a) / d = rd / d <<---- divide both sides by d (C-a) / d = r <<<--- simplify, since d/d = 1 Solved: r = (C-a)/d Does this make sense? For your next problem, when you post the question, go ahead and provide your best effort at solving for what the problem is asking,,, that will make it easier to make sure you understand or to help you learn if you are missing something. Good luck!!!!

OpenStudy (anonymous):

Thanks so much, it helped me a lot!

OpenStudy (anonymous):

I'm glad to help anyone who wants to learn :) Nice to know it was helpful for you :)